One of South Africa’s cellphone operators, Cell C, is changing their overall business strategy. The focus is now on providing customers with “affordable and accessible services." To do this they need your economic expertise.

They know that when price for premium services increases from R2/minute to R2.2/minute the demand for their services falls from 12 customers to 10 customers.

a. Using the point formula calculate the price elasticity of demand.

1
Expert's answer
2021-10-01T13:23:58-0400

Price elasticity of demand"=\frac{\%∆ Q}{\%∆ P}"

We calculate change in price and change in demand

Decrease in demand

"12-10=2\\\frac{2}{12}×100=17\%"

Increase in price

"2.2-2=0.2\\\frac{0.2}{2}×100=10\%"


Price elasticity of demand"=\frac{\%∆ Q}{\%∆ P}"

"\frac{17\%}{10\%}=1.7"